= 10v – bu, y = 2v – 5u implies a(u, v) 2(x,y)
Answer #1

Solution Given that х = 109 - вч = 29 - 5 и How 9 - 0 = 21 2 he - 1 11 24 - - 6 -7 24 ХК 6 Х. ка Q-Q) — ве re Э х ке. (о xlow